### v3

# Advancing Large Language Models for Tibetan with Curated Data and Continual Pre-Training

Abstract
Large language models have achieved remarkable progress across many languages. However, Tibetan, as a representative low-resource language, is particularly underrepresented in existing models due to the scarcity of high-quality training corpora. To address this gap, we curate the largest Tibetan pre-training corpus to date, aggregating data from diverse sources and applying a dedicated data cleaning and processing pipeline tailored for Tibetan. With the curated data, we continue pre/post-training a multilingual base model to enhance its generative capabilities in Tibetan. To evaluate the Tibetan capabilities of the model, we create new high-quality Tibetan benchmarks, and complement them with existing public benchmarks. Experimental results demonstrate that our model consistently and significantly outperforms both open-source models of similar scale and Tibetan-tailored models across a wide range of tasks.
Advancing Large Language Models for Tibetan with Curated Data and Continual Pre-Training
Leiyu Pan1, Bojian Xiong1, Lei Yang1, Renren Jin1,2, Shaowei Zhang1 Yue Chen1, Ling Shi1, Jiang Zhou1, Junru Wu1, Zhen Wang1, Jianxiang Peng1 Juesi Xiao1, Tianyu Dong1, Zhuowen Han1, Zhuo Chen1, Yuqi Ren1,2, Deyi Xiong1,2111Corresponding author. [^2] Project Lead. 1TJUNLP Lab, Tianjin University 2Yuanhui AI Lab
1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress, demonstrating exceptional performance across a wide range of natural language processing tasks Yang et al. (2025); Liu et al. (2024); Grattafiori et al. (2024), largely due to access to massive, high-quality training data Kaplan et al. (2020); Wettig et al. (2024). However, the pre-training corpora of mainstream LLMs are predominantly composed of high-resource languages such as English and Chinese Touvron et al. (2023); Grattafiori et al. (2024). This imbalance in data distribution has led to a strong performance bias toward high-resource languages, while support for low-resource languages remains significantly limited Zhu et al. (2024b); Sun et al. (2024).
The root causes of this disparity lie in the inherent scarcity of accessible data for low-resource languages, as well as the considerable challenges involved in data collection, cleaning, and standardization Kudugunta et al. (2023); Penedo et al. (2024a). These limitations hinder the development and effectiveness of LLMs for underrepresented languages, thereby preventing speakers of those languages from benefiting equally from the latest advances in artificial intelligence Nguyen et al. (2024b); Dou et al. (2024). Moreover, this technological imbalance risks further concentrating linguistic resources and innovation around dominant languages, posing a serious threat to global linguistic diversity Üstün et al. (2024) and cultural preservation Xu et al. (2024); Li et al. (2024). Therefore, advancing research on LLMs for low-resource languages is not only a matter of technological development but also a critical step toward promoting AI equity and safeguarding cultural diversity Shen et al. (2024).
Tibetan, as a representative low-resource language, is one of China’s minority languages and is spoken by millions of native speakers. It serves as a vital vehicle for the preservation of Tibetan culture and historical heritage Liang et al. (2024); Huang et al. (2025). Despite its significant population base and cultural value, Tibetan remains severely underrepresented in terms of digitized linguistic resources and has long been categorized as a low-resource language Liu et al. (2022); Liu and Best (2025). In recent years, the development of multilingual LLMs has led to gradual improvements in support for Tibetan Liu et al. (2024); Lv et al. (2025). However, due to the dominance of high-resource languages in training corpora and the limited scale, quality, and standardization of Tibetan data, existing models still struggle with tasks involving Tibetan understanding and generation Wenhao et al. (2024).
In light of the aforementioned challenges, we create a multilingual LLM supporting Tibetan, Chinese, and English. We present a comprehensive pipeline for building a Tibetan-capable LLM, covering data curation, pre-training, post-training, and evaluation.
To curate pre-training data for Tibetan, we build our Tibetan pre-training corpus from four primary sources: (1) publicly available datasets, (2) large-scale web crawling, (3) synthetic data generation, (4) proprietary data. To ensure high data quality, we also develop a customized data cleaning pipeline tailored for Tibetan. In total, we have collected 72 GB of cleaned Tibetan text, which, to the best of our knowledge, constitutes the largest Tibetan corpus ever constructed for LLM pre-training.
To build our model, we continue to pre-train Qwen2.5-7B-base model Qwen et al. (2025), a strong open-source multilingual LLM, to extend its capabilities to Tibetan. To facilitate effective cross-lingual transfer and accelerate the model’s adaptation, we incorporate both Chinese and English data alongside the curated Tibetan data during this continual pre-training phase. Following pre-training, we perform instruction tuning using multilingual instruction-following datasets collected from open-source repositories in Tibetan, Chinese and English, further enhancing ability of the model to handle diverse downstream tasks across the three languages.
For evaluation, given the limited availability of standardized Tibetan benchmarks for large language models, we construct multiple new Tibetan evaluation datasets by automatically translating existing benchmarks into Tibetan. All translations are rigorously proofread by native Tibetan speakers. The translated benchmarks are supplemented with existing public Tibetan benchmarks to provide a more comprehensive evaluation framework. Results show that both our base and chat models consistently outperform other models of similar scale across a variety of tasks, and in some cases, achieve performance comparable to significantly larger LLMs. Additionally, by extending the model vocabulary to include frequent Tibetan character combinations, we further enhance both generation quality and decoding efficiency, enabling faster inference under identical runtime conditions.
Our main contributions can be summarized as follows:
  • We enhance the foundation for Tibetan natural language processing by constructing a high-quality Tibetan pre-training corpus (the largest Tibetan dataset to date) and developing several evaluation benchmarks specifically designed for Tibetan LLMs.
  • With the curated data, we build a multilingual LLM with strong capabilities in Tibetan.
  • Extensive evaluations show that our model achieves strong improvements in accuracy compared to other open-source models and Tibetan-tailored models of similar scale across various benchmarks.
2 Related Work
In recent years, multilingual LLMs have made significant progress, gradually evolving from supporting a limited number of high-resource languages to covering a broader range of low- and medium-resource languages Yang et al. (2025). This improvement is largely driven by the continued expansion of multilingual corpora, both in terms of the number of supported languages and the amount of data available for each language Penedo et al. (2024b). As a result, the multilingual capabilities of LLMs have seen steady improvements Üstün et al. (2024); Ji et al. (2024). However, despite this progress, current multilingual models still provide insufficient support for Tibetan Qwen et al. (2025). Existing open-source multilingual corpora include very limited Tibetan data in terms of both scale and diversity, which significantly hampers model performance on Tibetan language understanding and generation Kudugunta et al. (2023); Nguyen et al. (2024a). To address this gap, we focus on building a large-scale, high-quality Tibetan pre-training corpus and training a dedicated language model with enhanced support for Tibetan, thus improving performance in downstream Tibetan tasks.
To efficiently leverage language-specific data for building models tailored to a particular language, a straightforward approach is to combine the collected data with multilingual corpora and train a new model from scratch Ji et al. (2024). However, this method is computationally expensive. As a result, prior research has proposed several low-cost alternatives to enhance existing models’ performance on specific languages. One line of work explores in-context learning techniques Cahyawijaya et al. (2024). While these methods are relatively efficient, they do not fundamentally improve the model’s inherent capabilities in the target language and typically assume that the base model already has a reasonable degree of support for that language. This limitation makes such methods less effective for truly low-resource languages like Tibetan, where existing models perform poorly to begin with. Another line of work adopts continued pre-training to adapt pre-trained models to underrepresented languages. Notable examples include Sailor Dou et al. (2024) and LLaMaTurk Toraman (2024), which demonstrate that extending multilingual models with targeted continual pre-training can effectively enhance their capabilities in specific languages. Inspired by this approach, we employ a continued pre-training approach to enhance the Tibetan language support of the Qwen2.5 model.
Several large language models with support for Tibetan understanding and generation have been proposed recently , e.g., TiLamb Wenhao et al. (2024), Sunshine Huang et al. (2025), and T-LLaMa Lv et al. (2025). Some of these models adopt parameter-efficient fine-tuning techniques like LoRA for continued pre-training Hu et al. (2022). In contrast, our approach uses full-parameter fine-tuning, offering greater flexibility and potential for adaptation. Moreover, compared to the aforementioned models, our Tibetan training corpus is larger in scale, richer in source diversity, and covers a wider range of linguistic scenarios. As a result, the model we trained on this corpus achieves superior performance on both our curated Tibetan benchmarks and existing public evaluation datasets.
3 Pre-training Data Curation
During the pre-training data curation phase, we collect Tibetan data from a wide range of sources. In addition to gathering all available open-source Tibetan corpora, we also include web-scraped content, synthetically generated data, and a portion of proprietary resources. To ensure high data quality, we design and apply a dedicated data cleaning pipeline specifically for Tibetan, which is used to process the entire dataset.
3.1 Data Collection
Our data collection covers four main sources: open-source data collection, web data crawling, synthetic data generation, and private data usage. The composition of these sources is detailed in Appendix A.
Open-Source Data Collection
We extensively collect Tibetan data from various open-source multilingual datasets, as well as from monolingual Tibetan dataset. Detailed information about the collected datasets is provided in Appendix B. The majority of the data originates from web sources, but it also includes content from other domains such as news, books, and Wikipedia, contributing to the overall diversity of the corpus.
Web Data Crawling
Although the open-source data we have collected includes a wealth of web-based content, it suffers from limited timeliness, with many recent materials not reflected in existing datasets. To address this limitation, we utilize the crawl4ai tool [^1] https://github.com/unclecode/crawl4ai to collect data from 55 Tibetan-language websites, resulting in a total of 204,125 webpages. The crawled data spans a wide range of domains, which are primarily on news, but also include content from books, Wikipedia, and other categories, contributing to the diversity of our Tibetan corpus. Detailed information of web data crawling is given in Appendix C.
Synthetic Data Generation
Synthetic data is often used to compensate for the lack of real-world data in low-resource settings, and it has been shown to effectively improve the performance of LLMs Abdin et al. (2024). Since Tibetan is a low-resource language with limited high-quality data, we employ automatic data synthesis to supplement our Tibetan corpus. Given that existed LLMs exhibit limited ability in Tibetan monolingual generation, we choose to generate Tibetan data via translation. To further promote cross-lingual transfer, we construct parallel corpora using both source language sentences and their Tibetan translations for training. Specifically, we sample a portion of the FineWeb-Edu Penedo et al. (2024a) and Cosmopedia Ben Allal et al. (2024) pre-training datasets, translate them into Tibetan using the Google Translate API, and organize the resulting parallel data using the instruction templates introduced by Zhu et al. (2024a).
Private Data Usage
In addition to publicly available data, we also utilize a set of private Tibetan-language digital resources, primarily consisting of books and classical materials. These resources are obtained from non-public but legally accessible sources, such as internal digital archives and institutional collaborations. All data used in this category contain no personally identifiable information or user-generated content, and are solely used for research and model training purposes.
3.2 Data Cleaning and Deduplication
After collecting the Tibetan corpus, we perform a comprehensive data cleaning process to ensure high data quality suitable for LLM training. Due to the unique characteristics of Tibetan script, we develop a customized data cleaning pipeline tailored to Tibetan, inspired by the FineWeb data processing workflow Penedo et al. (2024a). Our cleaning pipeline consists of two main stages: language filtering and quality filtering. For language filtering, we use the FastText language classifier Grave et al. (2018) to identify Tibetan text, removing any documents with a Tibetan confidence score below 0.5. For quality filtering, we apply a combination of five filtering strategies: Gopher duplicate filtering, Gopher quality filtering Rae et al. (2021), C4 quality filtering Raffel et al. (2020), FineWeb quality filtering and sensitive content filtering. Some of the heuristic rules in these procedures are slightly modified to better accommodate the characteristics of Tibetan text. Detailed descriptions of quality filtering are provided in Appendix D.
After completing the data cleaning process, we apply MinHash-based deduplication using the DataTrove tool [^2] https://github.com/huggingface/datatrove to remove highly redundant text segments. This step helps enhance the diversity and effectiveness of the pretraining data.
3.3 Data Composition
After cleaning and deduplicating the Tibetan data, we further enhance the model’s Tibetan capabilities by incorporating high-quality, multi-domain Chinese and English data Chen et al. (2024) into the continued pre-training process. This approach leverages cross-lingual generalization to transfer the model’s existing strengths in Chinese and English to Tibetan. Additionally, it functions as a data replay strategy, helping to prevent catastrophic forgetting of general knowledge that could arise if the model were trained exclusively on Tibetan data. Additionally, we adopt a balanced multilingual data ratio, aiming to maximize the model’s Tibetan performance by mitigating interference from high-resource languages while still benefiting from cross-lingual generalization. The data volumes used for continued pre-training in each language are summarized in Table 1.
Language Category Volume Token
English Books 9.8 G 2.77 B
Code 9.7 G 4.42 B
Encyclopedia 3 G 0.77 B
Mathematical Corpora 4.6 G 1.27 B
Academic Papers 5.8 G 1.47 B
QA Forums 2.7 G 0.90 B
Web Pages 13.9 G 3.35 B
Chinese Books 8.5 G 2.72 B
Encyclopedia 3.3 G 1.00 B
QA Forums 3 G 0.88 B
Web Pages 35.1 G 10.23 B
Tibetan Books 7.5 G 0.88 B
Web Pages 64.8 G 7.61 B
Table 1: Pre-training data volume for three languages.
4 Continual Pre-training
We select Qwen2.5-7B-base Qwen et al. (2025) as our foundation model due to its strong multilingual capabilities, which provides a solid foundation for cross-lingual transfer to low-resource languages such as Tibetan. To enhance its ability to process Tibetan text, we first expand the vocabulary to include commonly used Tibetan character combinations, enabling more efficient and accurate tokenization. We then adopt a two-stage continued pre-training strategy: a language-balanced pre-training stage with a short context window to transfer Chinese and English knowledge to Tibetan, followed by a long-context pre-training stage to extend the model’s context length. In future work, we plan to adopt the recently released Qwen3-8B-base model Yang et al. (2025) as our new foundation model, leveraging its improved multilingual performance to further enhance Tibetan language modeling.
4.1 Vocabulary Expansion
The original Qwen2.5 tokenizer tends to over-segment Tibetan characters, resulting in higher computational costs and reduced generation controllability. To address this, we extend its vocabulary by training a 15K Byte-level BPE tokenizer on a representative Tibetan corpus and merging it with the original vocabulary, removing duplicates. The final vocabulary contains 165,428 tokens, enabling more efficient and accurate Tibetan tokenization.
4.2 Stage 1: Language Balanced Pre-training
The first stage of continued pre-training aims to build foundational Tibetan capabilities by leveraging both Tibetan monolingual data and cross-lingual transfer from high-resource languages. To ensure knowledge retention and avoid catastrophic forgetting, we use a balanced mix of Chinese, English, and Tibetan data during training.
After tokenization using our extended tokenizer, the training dataset comprises approximately 15B Chinese tokens, 15B English tokens, and 8.5B Tibetan tokens. It is worth noting that although the raw volume of Tibetan data is substantial, the final token count is lower due to the higher compression efficiency of our tokenizer, which is specifically adapted for Tibetan.
In this stage, we perform continual pre-training starting from the Qwen2.5-7B-base model, preserving its original configuration. To improve training efficiency, we use a relatively short context window of 4K tokens. Additional training details are provided in Appendix E.
4.3 Stage 2: Long Context Pre-training
To enhance the model’s capability for processing long inputs, we conduct long context pre-training based on the first stage model to extend the context window to 32K tokens. Since Qwen2.5-7B-base extensively pre-training on 32K long contexts Qwen et al. (2025), only a minimal amount of training is required to achieve our objective.
The training data consist of trilingual mixed text in Chinese (0.56B), English (2.5B), and Tibetan (1.5B, which is not utilized in the first stage), totaling 4.56B tokens. This contributes to improving the model’s generalization capability for Tibetan long contexts. Regarding training parameters, we maintain consistency with Qwen2.5-7B-base and the first stage pre-training. Specific data distribution and training parameters are detailed in Appendix F.
We evaluate the model’s performance in long context comprehension through needle-in-a-haystack evaluation using Claude as the evaluator. Due to the scarcity of Tibetan data, we construct a needle-in-a-haystack dataset for Tibetan text and employ recent news articles as needles to avoid the influence of the model’s inherent knowledge on the evaluation. Experimental results demonstrate that the model after long context pre-training significantly outperforms the model that completed only the first stage pre-training. Detailed experiments are provided in Appendix F.
5 Post Training
After continual pre-training, the model has acquired foundational capabilities in Tibetan. To further enhance its ability to follow instructions, we perform instruction tuning using a diverse and richly curated set of instruction data. In terms of domain coverage, the open-source instruction data we have collected spans a wide range of areas, including instruction following, multi-turn dialogue, subject-specific knowledge, logical reasoning, translation, summarization and etc. This diverse coverage is designed to comprehensively enhance the model’s alignment capabilities across different tasks. From a linguistic perspective, to facilitate the transfer of instruction-following capabilities from Chinese and English to Tibetan, we construct a multilingual instruction dataset that covers three languages. Detailed information on the post-training data and process is provided in Appendix G.
6 Experiments
Due to the current lack of standardized Tibetan benchmarks for LLMs, we constructed four benchmarks to evaluate general Tibetan capabilities. Combined with two existing benchmarks for traditional Tibetan NLP tasks, this enables a comprehensive evaluation of model performance. Experimental results demonstrate that our model outperforms both open-source models of similar scale and Tibetan-tailored models across all evaluated tasks.
6.1 General Tibetan Benchmark Construction
Existing benchmarks for evaluating LLMs are predominantly designed for high-resource languages, with limited resources available for Tibetan. To facilitate systematic evaluation of Tibetan LLMs across a wide range of tasks, we constructed general-purpose Tibetan benchmarks based on the translation of several well-established English evaluation benchmarks, including HellaSwag Zellers et al. (2019), ARC Clark et al. (2018), Xcopa Ponti et al. (2020), and XStoryCloze Lin et al. (2021).
We leveraged the Claude-Sonnet-3.7 model, which demonstrates strong Tibetan generation capabilities, to translate the original datasets into Tibetan. To ensure the quality and reliability of the translated data, we performed a rigorous multi-stage review process covering semantic consistency, grammatical correctness, cultural appropriateness, and content safety. Detailed translation and quality review procedures are provided in Appendix H. Ultimately, we constructed four benchmarks: HellaSwag-bo, ARC-bo, Xcopa-bo, and XStoryCloze-bo.
6.2 Evaluation Benchmarks
We evaluated our model on a diverse set of Tibetan benchmarks, including our constructed tasks (HellaSwag-bo, ARC-bo, Xcopa-bo, XStoryCloze-bo) for reasoning and knowledge assessment, TibetanQA Sun et al. (2021) for reading comprehension, and FLORES-200 nll (2024) for multilingual translation. Detailed benchmark descriptions are provided in Appendix I.
Model Hellaswag-bo ARC-bo Xcope-bo Xstorycloze-bo TibetanQA
(acc) (acc_norm) (acc) (acc) (acc) (EM) (F1)
Ours-Base 30.30 36.20 44.32 59.80 60.80 49.43 66.15
Ours-Base-32k 30.09 36.58 44.62 58.60 60.60 38.42 55.52
Ours-SFT 30.64 36.71 48.39 57.80 61.86 46.15 63.15
Qwen3-8B 27.07 30.56 29.16 50.40 50.37 17.80 30.22
LLaMA3.1-8B-Instruct 27.28 32.00 29.31 51.60 51.69 36.09 53.04
Qwen2.5-7B-base 26.69 31.57 29.60 51.80 49.97 19.31 32.38
Qwen2.5-7B-Instruct 26.46 31.03 28.47 51.60 50.43 10.15 18.42
DeepSeek-R1-Distill-Llama-8B 25.96 30.67 30.38 50.40 48.50 9.88 17.99
Table 2: Performance comparison between our models and open-source models on different tasks.
Model Hellaswag-bo Arc-bo Xcope-bo Xstorycloze-bo TibetanQA Flores-200
(acc) (acc_norm) (acc) (acc) (acc) (EM) (F1) zh-bo en-bo
Ours-Base 30.30 36.20 44.32 59.80 60.80 49.43 66.15 13.81 9.20
Ours-Base-32k 30.09 36.58 44.62 58.60 60.60 38.42 55.52 20.86 13.56
Ours-SFT 30.64 36.71 48.39 57.80 61.86 46.15 63.15 21.56 11.42
Yak_Llama2_7B 24.67 26.40 25.64 53.00 48.37 0.12 0.24 11.81 11.10
Tibetan-Alpaca-7B 25.95 28.23 27.94 51.20 49.97 2.47 4.81 4.87 5.69
Tibetan-Llama2-7B 26.09 27.58 27.54 50.60 49.37 1.20 2.38 11.81 11.10
Table 3: Performance of our models compared to other Tibetan models on different benchmarks.
6.3 Comparison against Open-Source Models
We conducted a comparative evaluation of three versions of our model: Ours-Base, which is trained with one stage of continual pretraining; Ours-Base-32k, which undergoes two stages of continual pretraining; and Ours-SFT, which is further instruction-tuned. These models are evaluated against leading general-purpose open-source models on both standard reasoning and Tibetan-specific benchmarks.
On multiple-choice reasoning tasks such as Hellaswag-bo and ARC-bo, our models consistently outperform the baselines in accuracy scores. The advantage is even more pronounced on TibetanQA, where our models outperform general-purpose models by a large margin in both EM and F1 scores. These results underscore the effectiveness of our continued pretraining strategy, instruction tuning, and data construction pipeline in improving performance for low-resource languages like Tibetan.
6.4 Comparison against Existing Tibetan-Tailored Models
We further compared our models with several instruction-tuned models specifically designed for Tibetan. As shown in Table 3, our model consistently outperform existing Tibetan instruction-tuned models across all tasks. In particular, Ours-SFT achieves the strongest overall performance, showing substantial improvements in reading comprehension and question answering, where other models fall significantly behind. Moreover, in both zh-bo and en-bo translation tasks, our models achieve the highest BLEU scores among all evaluated models, demonstrating strong multilingual alignment and effective cross-lingual generalization. It is worth noting that our comparison only includes open-source Tibetan-tailored models. Models such as Sun-Shine Huang et al. (2025), which are not publicly available, are not included in this evaluation.
7 Conclusion
To address the limited support for Tibetan in existing models, we construct a high-quality, large-scale Tibetan pre-training corpus. With the curated corpus, we perform continual pretraining to develop a multilingual large language model supporting Tibetan, Chinese, and English. Our model demonstrates superior performance compared to both general-purpose open-source models and Tibetan-tailored models of similar scale on our constructed Tibetan benchmark as well as traditional Tibetan NLP tasks.
Limitations
Due to budget constraints, our current Tibetan pre-training corpus does not include data obtained through OCR of PDF documents. We are actively collecting Tibetan book PDFs to support future research in this direction. Additionally, limited computational resources have restricted our continued pre-training experiments to the 7B model scale. Exploring larger model scales will be left for future work.
Appendix A Tibetan Data Composition
To support large-scale Tibetan language modeling, we collect a diverse and high-quality training corpus from four major sources: open-source datasets, web data crawling, synthetic data generation, and private data. These data sources vary in style, domain coverage, and linguistic features, and together form a comprehensive Tibetan pre-training corpus. Table 4 summarizes the detail information of each data component.
Data Source Language Size (After Filter)
Open-source data Tibetan 9.53 G
Web data crawling Tibetan 4.18 G
Synthetic data Tibetan, Chinese, English 17.46 G
Private data Tibetan 41.13 G
Table 4: Overview of the four main sources of Tibetan training data.
Appendix B Open-Source Data Collection Details
The open-source Tibetan corpus we collected primarily comes from two sources: (1) the Tibetan portions of multilingual open-source datasets, and (2) dedicated monolingual Tibetan datasets. These corpora cover a wide range of domains. However, we observe that a portion of the web-based data originates from different snapshots of Common Crawl, which may introduce redundancy. To address this, we appliy deduplication to remove duplicate content across time.
Appendix C Web Data Crawling Details
We manually survey 45 Tibetan-language websites and collect web content from them. Most of these sites are news-oriented, and our primary goal is to obtain up-to-date Tibetan text with strong temporal relevance. During the web crawling process, we extract hyperlink elements from each page and determine whether the linked page belonged to the same root domain. If so, and the page has not been previously visited, we recursively crawl the subpage. Throughout the entire process, all crawling behavior strictly adhere to the rules defined in each site’s robots.txt file.
Appendix D Quality Filter Details
Gopher Duplicate Filtering
This method removes both documents with a large number of short, repeated segments and those with low duplication but excessively long content. We measure duplication at both the line and paragraph levels by computing the proportion of duplicated content and the proportion of characters within duplicated segments.
  • Character proportion in repeated n-grams: For 2-gram, 3-gram, and 4-gram patterns, we compute the proportion of characters contained in the most frequent repeated n-gram relative to the total number of characters in the document, with filtering thresholds set at 0.20, 0.18, and 0.16, respectively. In contrast, for n-grams ranging from 5 to 10, we calculate the proportion of characters contained in all repeated n-grams relative to the total character count. The corresponding thresholds decrease gradually from 0.15 for 5-grams to 0.10 for 10-grams.
  • Character proportion in duplicated lines: Documents in which duplicated lines contain more than 20% of the total characters are filtered out.
  • Duplication ratio: Documents are filtered out if more than 30% of their lines or 30% of their paragraphs are exact duplicates.
Gopher Quality Filtering
This stage filters out any documents that violate the quality heuristics defined in the Gopher data pipeline. Specifically, a document is removed if it meets any of the following conditions:
  • Contains fewer than 50 words or more than 10,000 words.
  • Has an average word length less than 2 or greater than 10.
  • Has a symbol-to-word ratio greater than 0.1.
  • Fewer than 80% of the words contain at least one alphabetic character.
  • More than 90% of the lines begin with bullet points.
  • More than 30% of the lines end with ellipses.
To better handle Tibetan text, we replaced the default NLTK tokenizer [^3] https://github.com/nltk/nltk with a custom tokenizer designed specifically for Tibetan. This allows for more accurate tokenization and filtering tailored to the unique linguistic features of Tibetan.
C4 Quality Filtering
The following filtering rules are applied to ensure data cleanliness, in line with the C4 dataset preprocessing strategy:
  • Only lines containing at least three words are retained.
  • Any page containing words from a predefined list of offensive, profane, obscene, or otherwise inappropriate terms is discarded.
  • All lines containing the word "Javascript" are removed.
  • Any page containing the phrase "lorem ipsum" is discarded.
  • Any page containing curly braces is removed.
  • Since some crawled pages originated from Wikipedia and included citation markers, all pages containing such citation patterns are excluded.
  • To eliminate templated or boilerplate legal content, we remove any line that contains phrases such as “terms of use”, “privacy policy”, “cookie policy”, “uses cookies”, “use of cookie”, or “use cookie”.
FineWeb Quality Filtering
Documents that violate the FineWeb quality heuristics are removed. Specifically, a document is filtered out if it meets any of the following conditions:
  • More than 67% of its lines have a length less than or equal to 30 characters.
  • After removing empty lines, the proportion of repeated characters exceeds 1% of the total character count.
  • The ratio of newline characters to total word count exceeds 0.3.
Sensitive Content Filtering
To eliminate politically sensitive content from the collected data, we manually curated a blacklist of sensitive terms. Any document containing one or more terms from this list is filtered out.
Appendix E Continual Pre-training Details
To improve the training efficiency of the the model, we adopt a packing strategy Raffel et al. (2020) to preprocess the data. This involves concatenating and truncating the data to ensure that each training sample is exactly 4096 tokens in length. Given the large volume of training data, we also apply a pre-tokenization strategy, where the tokenized and packed datasets are preprocessed and saved in advance to streamline the subsequent training process.
For the training framework, we use the LLaMA-Factory library Zheng et al. (2024) in combination with DeepSpeed, enabling the ZeRO Stage 2 optimization strategy Rajbhandari et al. (2020). Training is conducted using 32 A800 GPUs with bf16 precision. Regarding optimization, we use the standard AdamW optimizer Loshchilov and Hutter (2017) with a maximum gradient norm of 1.0. The learning rate is set to 5e-5, with a warmup ratio of 0.1, and we adopt a cosine learning rate scheduler.
Appendix F Long Context Pre-training Details
Data Distribution. To fully account for the distributional characteristics of the training data in the first stage and further enhance the model’s long context processing capabilities in Tibetan, we introduce longer mixed-language texts (Chinese, English, and Tibetan) as training corpora in the second training stage.We sample 0.56B tokens of Chinese data, 2.5B tokens of English data, and 1.5B tokens of Tibetan data (unused in the first stage) from long context sources in each respective language, resulting in a trilingual long context training set totaling 4.56B tokens.
Training Parameters. For training configurations, we maintain consistency with the Qwen2.5-7B-base model by setting the RoPE $\theta$ parameter to 1,000,000. The global batch size is set to 128 to ensure consistency in the number of tokens per batch with the first stage pre-training. Additionally, the learning rate is reduced from 5e-5 to 1e-5, while all other hyperparameters remained unchanged. This stage of training is conducted using DeepSpeed ZeRO-3 Offload strategy for a total of 1,000 steps.
Experimental Results. We evaluate the performance of four model variants on the needle-in-a-haystack task: the base and SFT models from the first stage, and the base and SFT models after long context training in the second stage. The evaluation is conducted using bilingual long context samples in English and Tibetan. As illustrated in Figure 1, the results demonstrate that long context training successfully extend the model’s context window to 32K tokens and significantly improve its multilingual capabilities.
(a) 4K_base/en
(b) 4K_base/bo
(c) 4K_SFT/en
(d) 4K_SFT/bo
(e) 32K_base/en
(f) 32K_base/bo
(g) 32K_SFT/en
(h) 32K_SFT/bo
Figure 1: Performance of the first stage 4K context window base and SFT models, and the second stage long context training 32K context window base and SFT models on the English (en) and Tibetan (bo) needle-in-a-haystack task.
Appendix G Post Training Details
Training Data
We extensively collect open-source instruction datasets in Chinese, English, and Tibetan. Upon analysis, we find that most existing Tibetan instruction datasets are generated by translating open-source Chinese and English instruction data using Google Translate. While this approach is simple and convenient, it often results in low-quality Tibetan instructions due to the current limitations of Google’s Tibetan translation capabilities. Additionally, such translations may introduce unintended biases. To address these issues, we collaborate with native Tibetan speakers to produce additional high-quality translation and summarization data. These samples are organized using instruction-style templates to further enhance the model’s instruction-following abilities in Tibetan, particularly in the areas of translation and summarization.
Training Setting
We apply a packing strategy to process the instruction data and adopt the standard supervised fine-tuning loss computation method. Specifically, the loss is calculated only on assistant tokens, while system prompt and user tokens are excluded. The learning rate is set to 1e-5, the maximum gradient norm was set to 1.0, and the model is trained for three epochs.
Appendix H General Tibetan Benchmark Construction Details
To construct general-purpose Tibetan benchmarks for evaluating LLMs, we translate several widely used English evaluation datasets (e.g., HellaSwag, ARC, XCOPA, and XStoryCloze) into Tibetan. These datasets span a range of reasoning, commonsense, and knowledge-intensive tasks, making them suitable for comprehensive capability assessment.
Translation
We use the Claude-Sonnet-3.7 model to perform English-to-Tibetan translation. This model is equipped with advanced capabilities for Tibetan generation, enabled by large-scale cross-lingual pre-training.
Quality Review
To ensure translation quality, we perform a structured manual review process guided by our internal Data Review Guidelines. The review consists of three main stages:
  • Semantic and Terminological Accuracy: Reviewers compare the source and translated texts to identify semantic deviations, including incorrect interpretation of core concepts, logical inconsistencies, or mistranslation of technical terms.
  • Content Compliance: All translated content is checked for potentially harmful material, such as violent or discriminatory language. Additionally, reviewers filter out content that may conflict with Tibetan cultural norms or sensitivities.
  • Linguistic and Formatting Correctness: Reviewers examine grammatical structures (e.g., use of case particles, honorifics), fluency, and formatting consistency, including punctuation, line breaks, and symbol pairing.
Translations exhibiting conspicuous machine translation artifacts or unnatural expressions shall be rejected under the "Low-Quality Machine Translation" category to ensure training data aligns with Tibetan linguistic conventions. During review, annotators must maintain objective assessment, comprehensively label rejection categories for multi-issue samples with specific problem descriptions, thereby constructing a high-quality Tibetan evaluation benchmarks.
Appendix I Evaluation Benchmarks Details
The evaluation covers six key benchmarks, each designed to assess a specific capability of the model within the Tibetan language setting:
  • HellaSwag-bo: Evaluates commonsense reasoning about event progression using multiple-choice questions. Accuracy is measured against adversarial distractors to test the model’s ability to make logically coherent predictions across diverse scenarios.
  • Xcopa-bo: Focuses on causal reasoning by requiring the model to choose the correct cause or effect of a given event.
  • XstoryCloze-bo: Assesses zero-shot narrative understanding by requiring the model to choose the most plausible story ending. It evaluates both temporal consistency and semantic coherence in Tibetan.
  • ARC-bo: Benchmarks factual knowledge by evaluating middle-school level science questions that require basic recall. Accuracy reflects the model’s ability to retrieve simple facts in Tibetan.
  • TibetanQA: A large-scale machine reading comprehension dataset containing 20,000 QA pairs across 12 domains and four difficulty levels: word matching, synonym substitution, multi-sentence reasoning, and ambiguous questions. It tests fine-grained Tibetan text understanding.
  • FLORES-200 (Tibetan subset): Used to evaluate translation quality among Chinese, English, and Tibetan. The dataset consists of professionally verified parallel sentences and tests character handling, grammatical correctness, and translation fidelity.
Appendix J Tokenizer compression ratio evaluation
In the tokenizer compression ratio evaluation, our Tibetan model demonstrates exceptional performance, achieving a compression ratio of 3.9644. This figure far surpasses those of mainstream models in the industry—specifically, Qwen3-8B’s 0.7315 and Llama3-8B’s 0.4885—highlighting a significant advantage in compression capability. A high compression ratio means that when processing Tibetan texts of the same length, the model requires fewer tokens, which effectively reduces computational load and memory usage, thereby significantly accelerating the inference process. This feature not only enhances the operational efficiency of the model but also provides strong support for the application of Tibetan natural language processing tasks in real-time response scenarios and resource-constrained environments, underscoring the model’s prominent value in optimizing inference speed and reducing computational costs.